OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

−3(x+7)=−12 Step 1: Simplify both sides of the equation. −3x−21=−12 Step 2: Add 21 to both sides. −3x−21+21=−12+21 −3x=9 Step 3: Divide both sides by -3. −3x −3 = 9 −3 x=−3

OpenStudy (anonymous):

x = 5 over 3 , x = − 19 over 3 x = −3, x = −11 x = −3, x = 11 No solution these are the options so would it be the 2nd one?